Hence find these curvatures in the following cases:
i. f(u) = e^u, g(u) = u.
ii. f(u) = 2 + sin u, g(u) = u.
1
Expert's answer
2020-11-18T19:06:49-0500

"k=\frac{|[\overline{r}^\prime,\overline{r}^{\prime\prime}]|}{|\overline{r}^\prime|^3}\\
i.\ x=u, y=e^u, z=0\\
\overline{r}(u)=(u,e^u,0)\\
\overline{r}^\prime(u)=(1,e^u,0)\\
\overline{r}^{\prime\prime}(u)=(0,e^u,0)\\
[\overline{r}^\prime,\overline{r}^{\prime\prime}]=(0,0,e^u)\text{ (vector product of } \overline{r}^\prime\text{ and }\overline{r}^{\prime\prime}).\\
|[\overline{r}^\prime,\overline{r}^{\prime\prime}]|=e^u\\
|\overline{r}^\prime(u)|=\sqrt{1+e^{2u}}\\
k=\frac{e^u}{(\sqrt{1+e^{2u}})^3}\\
ii.\ \ x=u, y=2+\sin u, z=0\\
\overline{r}(u)=(u,2+\sin u,0)\\
\overline{r}^\prime(u)=(1,\cos u,0)\\
\overline{r}^{\prime\prime}(u)=(0,-\sin u,0)\\
[\overline{r}^\prime,\overline{r}^{\prime\prime}]=(0,0,-\sin u)\text{ (vector product of } \overline{r}^\prime\text{ and }\overline{r}^{\prime\prime}).\\
|[\overline{r}^\prime,\overline{r}^{\prime\prime}]|=|\sin u|\\
|\overline{r}^\prime(u)|=\sqrt{1+\cos^2u}\\
k=\frac{|\sin u|}{(\sqrt{1+\cos^2u})^3}\\"